OpenStudy (anonymous):

What is the sign of the product (–7)(–2)(–5)(1)? Positive, because the products (–7)(–2) and (–5)(1) are negative, and the product of two negative numbers is positive Positive, because the products (–7)(–2) and (–5)(1) are positive, and the product of two positive numbers is positive Negative, because the product (–7)(–2) is positive, and the product (–5)(1) is negative and the product of a positive and a negative number is negative Negative, because the products (–7) (–2) and (–5) (1) are negative and the product of two negative numbers is negative

OpenStudy (anonymous):

@Jhannybean @e.mccormick

OpenStudy (anonymous):

@AndrewTheCookie @Michele_Laino

OpenStudy (anonymous):

Pls i just need the answers. I know how to solve them but im in a hurry

OpenStudy (michele_laino):

please note that: \[(-)*(-)=(+)\] so: \[(-)*(-)*(-)=(+)*(-)=(-)\]

OpenStudy (michele_laino):

so?

OpenStudy (anonymous):

I'm confused between A and B

OpenStudy (michele_laino):

I think that they are both wrong!

OpenStudy (anonymous):

Really?

OpenStudy (michele_laino):

really!

OpenStudy (anonymous):

So then which one is it??

OpenStudy (michele_laino):

it is C.

OpenStudy (anonymous):

Thanks can you help with some more??

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Eva is showing her work in simplifying (7.3 – 7.2) – 3.8 + 3.7. In which step did Eva make an error? Step 1: (7.3 – 7.2) – 3.8 + 3.7 Step 2: 7.3 + (–7.2 – 3.8) + 3.7 (commutative property) Step 3: 7.3 – 11 + 3.7 Step 4: 7.3 + 3.7 – 11 (commutative property) Step 5: 11 – 11 = 0 Step 2; she wrote commutative instead of distributive Step 4; she wrote commutative instead of distributive Step 4; she wrote commutative instead of associative Step 2; she wrote commutative instead of associative

OpenStudy (michele_laino):

I think it is: "Step 2; she wrote commutative instead of associative"

OpenStudy (anonymous):

Thanks@!!!

OpenStudy (michele_laino):

because at step 2, Eva has applied the associative property of addition

OpenStudy (anonymous):

Simplify Negative 3 over 2 ÷ 9 over 6. –4 –1 4 1

OpenStudy (anonymous):

divided sign then -

OpenStudy (michele_laino):

yes I think, namely: \[\frac{ \frac{ -3 }{ 2 } }{ \frac{ 9 }{ 6 } }=\] the first fraction times the inverse of the second fraction, namely: \[\frac{ -3 }{ 2 }*\frac{ 6 }{ 9 }=...\]

OpenStudy (michele_laino):

so?

OpenStudy (anonymous):

-18 / 18

OpenStudy (michele_laino):

and then?

OpenStudy (anonymous):

divide?

OpenStudy (michele_laino):

\[\frac{ 18 }{ 18 }=1\]

OpenStudy (anonymous):

d?

OpenStudy (michele_laino):

so: \[\frac{ -18 }{ 18 }=...\]

OpenStudy (anonymous):

-1

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

so B.

OpenStudy (anonymous):

Tanisha opened a bank account. She deposited $74.50 into her account every month for 10 months. She used $32.50 every month to pay for painting classes. After 10 months, she used 1 over 2 of the total money left in her account to go to a summer camp for painters. What is the total amount of money Tanisha spent to go to the summer camp? $210.00 $325.00 $420.00 $840.00

OpenStudy (anonymous):

A 210.000

OpenStudy (anonymous):

$210.00

OpenStudy (michele_laino):

every mont Tanisha deposits 74.50$, and every month she spent 32.50$, so every month Tanisha deposit is equal to 42.00$. Now after 10 months, Tanisha bank account is of 42*10=420.00$ Since Tanisha uses only one half of that sum, she uses 420/2=...

OpenStudy (anonymous):

Simplify negative 2 and 1 over 9 – negative 4 and 1 over 3. 2 and 2 over 9 negative 6 and 4 over 9 negative 2 and 2 over 9 6 and 4 over 9

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

the ? = -

OpenStudy (michele_laino):

2 and 1/9=19/9

OpenStudy (michele_laino):

negative 4 and 1 over 3=-(4 1/3)=-13/3 so you have to calculate this: \[\frac{ 19 }{ 9 }-\left( -\frac{ 4 }{ 3 } \right)=...\]

OpenStudy (anonymous):

3.4

OpenStudy (anonymous):

no, 3.5 @Michele_Laino

OpenStudy (michele_laino):

oops sorry: so you have to calculate this: \[\frac{ 19 }{ 9 }-\left( -\frac{ 13 }{ 3 } \right)\]

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

6.5 @Michele_Laino

OpenStudy (michele_laino):

please, note that: \[\frac{ 19 }{ 9 }-\left( -\frac{ 13 }{ 3 } \right)=\frac{ 19 }{ 9 }+\frac{ 13 }{ 3 }=...\]

OpenStudy (anonymous):

It is 6.5

OpenStudy (anonymous):

32/12?

OpenStudy (michele_laino):

wait, please... \[\frac{ 19 }{ 9 }+\frac{ 26 }{ 9 }=\frac{ 45 }{ 9 }=5\]

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

You don't have to explain it though ;)

OpenStudy (michele_laino):

no, please I have made an error of sign, we have to calculate this:

OpenStudy (michele_laino):

\[-\frac{ 19 }{ 9 }-\left( -\frac{ 13 }{ 3 } \right)=-\frac{ 19 }{ 9 }+\frac{ 13 }{ 3 }=\] \[=-\frac{ 19 }{ 9 }+\frac{ 26 }{ 9 }=\frac{ 7 }{ 9 }\]

OpenStudy (michele_laino):

pleas another error, I'm very sorry: \[=-\frac{ 19 }{ 9 }+\frac{ 39 }{ 9 }=\frac{ 20 }{ 9 }=\] = 2 and 2/9

OpenStudy (michele_laino):

so A.

OpenStudy (michele_laino):

because: \[\frac{ 20 }{ 9 }=2+\frac{ 2 }{ 9 }\]

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Sally swam a total distance of 3 and 1 over 2 miles during the months of November and December. If Sally only swam 1 over 4 of a mile every day, which expression shows the number of days she went swimming? 3 and 1 over 2 ⋅ 1 over 4 3 and 1 over 2 ÷ 1 over 4 3 and 1 over 2 + 1 over 4 3 and 1 over 2 – 1 over 4

OpenStudy (anonymous):

What about this one @Michele_Laino

OpenStudy (michele_laino):

I think B. because we have to divide the total distance over the unitary distance, namely the distance that Sally swam every day